OpenStudy (anonymous):

Write (3y)^2 without exponents

OpenStudy (kc_kennylau):

how many times of \(y\) is \(3y\)? :)

OpenStudy (anonymous):

idk

OpenStudy (kc_kennylau):

\[(3y)^2\]\[=(3\times y)^2\]\[=(3\times y)(3\times y)\]\[=3\times y\times3\times y\]\[=9y^2\]

OpenStudy (anonymous):

without exponentssss

OpenStudy (kc_kennylau):

\[9y\times y\]

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

3y * 3y

OpenStudy (kc_kennylau):

so you have the answer